OpenStudy (anonymous):

Factor: 18ax-6ay+9bx-3by I got 3 as the GCF to get: 3(6ax-2ax+3bx-by)

OpenStudy (zzr0ck3r):

18ax-6ay+9bx-3by = 6a(3x-y) +3b(3x-y) = (6a+3b)(3x-y)= 3(2a+b)(3x-y)

OpenStudy (anonymous):

I did not see it that way, thank you. So for problems like that, you split the two in half to split the GCF?

OpenStudy (zzr0ck3r):

hmm, I dont think I understand what you are saying. we could have went your rout as well... 18ax-6ay+9bx-3by = 3(6ax-2ay+3bx-by) = 3(2a(3x-y)+b(3x-y)) = 3(2a+b)(3x-y)

OpenStudy (zzr0ck3r):

hope that helps

OpenStudy (anonymous):

It does, thank you! :)